OpenStudy (anonymous):

what is the inverse of f(x)=x^(2/3) for x greater than or equal to zero.

OpenStudy (anonymous):

I got y=x^(3/2), is that right?

OpenStudy (anonymous):

inverse is replacing x with y. you can rewrite it as y = x^(2/3) the inverse is simply x = y^(2/3) ----> y = x^(3/2)

OpenStudy (lalaly):

yeh thats 100% rightt! ..